You just signed a business consulting contract with one of your clients. The client will pay you $30,000 a year for 5 years for the service you will provide over this period. You anticipate the general inflation rate over this period to be 3.5%. If your desired inflation-free interest rate to be 14%, what is the worth of the 5th payment in today’s dollars?
Answer #1

Solution:-

To Calculate worth of 5th payment in today's dollar-

Present Value = \frac{Cash Flows * (1+Inflation)^{n}}{(1+r)^{n}}

Present Worth = \frac{30,000 * (1+0.035)^{4}}{(1+0.14)^{5}}

Present Worth = \frac{34,425.69}{1.92541458}

Present Worth = $17,879.62

If you have any query related to answer then please feel free to ask me in a comment. Thanks.